# Retrieval-Augmented Generation in Biomedicine: A Survey of Technologies, Datasets, and Clinical Applications

### v2

@@ Introduction
DRAGON-AI 83 OpenAI Text-Embedding 84 with Chroma DB Ontology generation 2024
RISE 85 OpenAI Text-Embedding with FAISS Diabetes-related inquiries 2024
Dual RAG 86 UPstage API and OpenAI Text-Embedding Diabetes management 2024
FAVOR-GPT 87 Open AI Text-Embedding with Weaviate DB Genome variant annotations 2024
Endo-chat 88 Open AI Text-Embedding with Faiss medical QA for gastrointestinal endoscopy 2024
ChatENT 89 Open AI Text-Embedding Medical QA in otolaryngology 2024
GuideGPT 90 Open AI Text-Embedding MRONJ QA on prevention, diagnosis, and treatment 2024
RECTIFIER 91 Open AI Text-Embedding with Faiss Clinical trial screening for heart failure patients 2024
RAP 92 Amazon Titan Text Embeddings v2 93 Nutrition-related question answering 2025
InfectA-Chat 94 Open AI Text-Embedding Infectious disease monitoring and QA 2025
DCRAG 95 OpenAI Text-Embedding Auto-annotation of plant phenotype 2025
@@
PM-Search 24 BioBERT 96 Clinical literature retrieval 2022
LADER 97 PubMedBERT 98 Biomedical literature retrieval 2023
CLEAR 99 BioBERT Clinical information extraction 2025
KG-RAG 100 PubMedBERT Biomedical Multiple-choice questions 2024
WeiseEule 101 MedCPT, BioBERT, BioGPT 102 with Pinecone DB Biomedical QA 2024
@@
BiomedRAG 4 MedLLaMA 13b Biomedical NLP tasks 2024
SurgeryLLM 28 GPT4All 103 with Chroma DB Surgical decision support 2024
Myers et al. 104 LLM2Vec (for comparison) 105 Clinical Information retrieval 2024
@@
Self-BioRAG 106 MedCPT 33 Medical QA 2024
@@
WeiseEule 101 MedCPT, BioBERT, BioGPT with Pinecone DB Biomedical QA 2024
CliniqIR 52 MedCPT and BM25 Diagnostic decision support 2024
@@
MEDRAG 101 BM25, SPECTER 22, Contriever 57and MedCPT Biomedical RAG Tool 2024
@@
RAMIE 107 MedCPT, Contriever and BMRetriever 55 Biomedical IR about dietary supplements 2025
@@
Type 2: Commercial embedding APIs. The proliferation of foundation models has led to the emergence of proprietary embedding solutions offered by major AI private organizations. These commercial embedding APIs have gained significant adoption in biomedical research, despite their associated costs, due to several compelling advantages: i) they typically leverage substantially larger pretraining datasets, ii) benefit from extensive computational resources during development, and iii) receive continuous refinements based on diverse user interactions. As shown in Table LABEL:tab:Classification, OpenAI’s embedding models dominate this category in recent studies, underscoring their enhanced semantic representation capabilities. While OpenAI’s offerings dominate, alternatives are emerging. The RAP system 92 utilizes Amazon’s Titan Text Embeddings v2 93 for nutrition-related question answering. Other commercial APIs such as Anthropic’s Voyage-2 108 and Google’s Vertex AI embeddings 109, though not specifically represented in Table LABEL:tab:Classification, are alternatives for biomedical RAG applications.
@@
Type 3: Domain adaptation PLMs. Domain-adapted PLMs are language models that have been specifically pre-trained or fine-tuned on biomedical and clinical corpora, allowing them to develop specialized representations of medical terminology, concepts, and relationships 96, 98. These models mitigate a fundamental challenge in biomedical information retrieval: general-domain language models often struggle with the highly specialized vocabulary, complex semantic relationships, and unique linguistic patterns prevalent in biomedical literature 110. By adapting pre-training to include PubMed articles, clinical notes, or other domain-specific corpus, these models develop a more nuanced understanding of biomedical language, resulting in significantly improved performance on retrieval tasks involving medical concepts and relationships compared to their general-domain counterparts 110. As shown in Table LABEL:tab:Classification, examples of these domain-adapted models include PubMedBERT 98, BioBERT 96 and BioGPT 102, which have been widely adopted in biomedical RAG systems. For example, PM-Search 24 utilized BioBERT for clinical literature retrieval, effectively leveraging the model’s domain-specific representations to improve search precision. Similarly, PubMedBERT has been successfully deployed in LADER 97 for biomedical literature retrieval and in KG-RAG 100 for answering biomedical multiple-choice questions, demonstrating its versatility across different tasks. More recent systems like WeiseEule 101 combine multiple biomedical models including BioBERT, BioGPT, and MedCPT 33 to achieve enhanced performance in biomedical QA.
@@
Type 4: LLM-based embedding. Generative LLMs have demonstrated remarkable generative capabilities across diverse NLP tasks, owing to their sophisticated understanding of language and context. With billions of parameters and training on vast corpora, these models develop rich internal representations that capture nuanced semantic relationships and contextual dependencies in text 12. Recognizing this powerful language comprehension ability, researchers have begun leveraging open-source LLMs as embedding generators for retrieval tasks, extracting representations from their internal layers to create semantically meaningful document and query embeddings that outperform traditional embedding approaches. As shown in Table LABEL:tab:Classification, notable implementations include BiomedRAG 4, which utilizes MedLLaMA 13b 111, SurgeryLLM 28, which employs GPT4All 103, and the comparative study by Myers et al. 104, which evaluates LLM2Vec 105 for clinical information retrieval applications. However, most LLMs are decoder-only architectures that employ causal attention mechanisms, which inherently limit their ability to generate rich contextualized representations. Thus, researchers are increasingly focusing on developing more advanced LLM-based embedding methods, such as Llama2Vec 112, Landmark Embedding 113, which provide promising dense retrieval alternatives in biomedical RAG.
@@
Type 5: Domain adaptation retrievers. Although the aforementioned off-the-shelf retrievers are readily available, the task of searching for relevant and accurate documents in the medical domain remains challenging. Consequently, customized retrievers tailored specifically to scientific and biomedical tasks have been developed. As shown in Table LABEL:tab:Classification, these advanced retrievers include Contriever 57, MedCPT 33, SPECTER 22, BMRetriever 55, and the MEDRAG toolkit 4. These specialized retrieval architectures demonstrate the importance of domain adaptation in biomedical information retrieval, significantly improving retrieval performance in several studies 106, 52, 107. For example, Jeong et al. 106 proposed Self-BioRAG, which leverages MedCPT to improve medical question answering through a self-augmentation approach. Similarly, the MEDRAG framework 101 combines multiple retrieval strategies, including BM25, SPECTER, Contriever, and MedCPT to create a comprehensive biomedical RAG tool. RAMIE utilizes MedCPT, Contriever, and BMRetriever as retrievers, builds a RAG framework to extract diverse types of information about dietary supplements (DSs) from clinical records, enhanced overall accuracy in its experiments 107. These implementations highlight how domain-adapted retrievers benefit biomedical RAG systems.
Dense Retriever Infrastructure. Vector stores provide an infrastructure element for dense retrievers in semantic search systems 114. These specialized databases efficiently index, store, and search high-dimensional vector embeddings, offering optimized methods like K-nearest neighbors (KNN) for RAG developers. Vector databases enable rapid retrieval of semantically similar documents at scale, significantly enhancing the performance and response time of biomedical RAG systems. The process of embedding, storing, and searching documents involves transforming text into vector representations, organizing them in specialized data structures, and efficiently retrieving the most relevant matches. Based on our survey of biomedical RAG publications presented in Table LABEL:tab:Classification, we have compiled the most commonly used vector databases in Table 3. In Table LABEL:tab:Classification, it can be observed that the most popular vector stores are FAISS and Chroma, which were utilized 4 times and 3 times respectively.
@@
Chroma 115 AI-native open-source vector database with built-in functionality
Faiss 116 Library for efficient similarity search and clustering of dense vectors
Pinecone 117 Production-ready vector database for similarity search at scale
Weaviate 118 Graph-based vector search engine with semantic capabilities
HNSW 78 Hierarchical navigable small world graphs for approximate nearest neighbor search
@@
Knowledge graph retrievers offer significant advantages in biomedical RAG by providing structured, interlinked domain knowledge that overcomes several limitations of traditional retrieval methods. Unlike vector-based approaches, knowledge graphs capture explicit relationships between biomedical entities, enabling more precise retrieval of contextually relevant information, enhanced reasoning capabilities, and improved explainability119. Table 4 presents knowledge graph-based biomedical RAG studies published between 2024 and 2025, outlining each system’s application domains, and knowledge graph source. As we can notice, the application areas vary significantly, including medical question answering 100, knowledge extraction 120, and diagnosis prediction 17.
@@
KG-RAG 100 Biomedical text generation, medical question answering SPOKE biomedical knowledge graph 2024
HEALIE 29 Personalized medical content generation HEALIE KG (customized) 2024
KRAGEN 121 Biomedical problem solving Alzheimer’s knowledge graph (AlzKB) 2024
Ascle 30 Various medical text generation tasks UMLS 2024
DALK 122 Alzheimer’s QA Alzheimer’s knowledge graph (AlzKB) 2024
Gilbert et al. 123 Medical information structuring and interlinking Not specified 2024
NetMe 2.0 120 Biomedical knowledge extraction BKGs constructed using OntoTagMe and Wikidata 2024
NEKO 124 Knowledge mining in synthetic biology PubMed-derived knowledge graphs 2025
DR.KNOWS 17 Diagnosis prediction from EHRs UMLS 2025
ESCARGOT 125 Biomedical reasoning and knowledge retrieval Alzheimer’s knowledge graph (AlzKB) 2025
As shown in Fig 2, the integration of knowledge graphs with LLMs in biomedical RAG applications can be analyzed through four key components:
Figure 2: The Framework of Knowledge Graphs Based Biomedical RAG
  • Biomedical knowledge graph construction or curation involves the systematic process of building, updating, or adapting knowledge graphs from domain-specific ontologies, literature, and other data sources to create structured biomedical knowledge repositories. NetMe 2.0 120 constructs biomedical knowledge graphs on-the-fly with OntoTagMe for entity linking while NEKO 124 mines PubMed literature for synthetic biology knowledge graphs. Differently, KG-RAG 100 utilizes the pre-existing SPOKE biomedical knowledge graph.
  • Query structuration encompasses the conversion of natural language queries or user inputs into structured queries that efficiently retrieve relevant subgraphs from the knowledge graph. ESCARGOT 125 employs Cypher queries for precise retrieval, KRAGEN 121 vectorizes knowledge graphs for semantic searching, and DR.KNOWS 17 transforms queries to retrieve UMLS-based diagnostic pathways aligned with patient data.
  • Path-based or entity-centric knowledge extraction refers to the process of extracting relevant knowledge patterns, paths, or entity-centered information from the knowledge graph based on the transformed query. DR.KNOWS 17 utilizes attention-based path ranking for diagnostic reasoning, KRAGEN 121 applies graph-of-thoughts for multi-hop relationships, and HEALIE 29 extracts paths incorporating both medical knowledge and socio-economic factors.
  • Strategic integration into LLM prompts involves methods for incorporating retrieved knowledge into prompts for LLMs in ways that maximize factual accuracy while optimizing token usage and computational efficiency. KG-RAG 100 implements token-optimized retrieval to reduce consumption, ESCARGOT 125 selectively integrates knowledge through dynamic Graph of Thoughts, and NetMe 2.0 120 transforms graph data into natural language prompts via its Graph-RAG module.
@@
Multimodal biomedical RAG represents an advanced approach to healthcare AI that integrates LLMs with multiple types of data inputs beyond text, such as medical images, electrocardiogram (ECG), and other clinical measurements. These systems can enhance diagnostic accuracy and clinical decision support by leveraging the complementary strengths of different data modalities, retrieving relevant information from specialized knowledge bases, and generating comprehensive insights that would be impossible to achieve with single-modality approaches 126. As shown in Table 5, multimodal biomedical RAG systems have been successfully deployed for amblyopia detection 32, prescription label interpretation 127, lung cancer staging 128, digital pathology analysis 129, and automated radiology report generation 130, 31, 131, 132, 133.
@@
Ranjit et al. 130 Chest X-Ray report generation Chest X-ray and text 2023
FactMM-RAG 31 Radiology report generation X-Ray and text 2024
MMed-RAG 131 Medical VQA & Report generation Medical images and text 2024
MEAG 32 Amblyopia diagnosis Eye tracking and text 2024
Tozuka et al.  128 Lung cancer staging CT findings and TNM classifications 2024
GPT4DFCI-RAG 129 Digital pathology QA Pathology images and text 2024
Raminedi et al. 19 Radiology report generation X-ray images and text 2024
RULE 132 Medical VQA & Report generation Medical images and text 2024
Thetbanthad et al. 127 Prescription label identification Image (labels) and text 2025
Hu et al. 134 Pathology report generation Whole slide images and text 2025
STREAM 133 Chest X-ray report generation Chest X-ray and text 2025
Figure 3 illustrates the architectural workflow of a multimodal biomedical RAG system. The key distinction between multimodal biomedical RAG and traditional biomedical RAG lies in data input diversity and processing architecture. While traditional RAG systems primarily process textual queries to retrieve text-based medical knowledge, multimodal RAG incorporates and processes non-textual data (images, sensor readings, etc.) alongside text, requiring specialized encoders for each modality to meaningfully integrate these diverse inputs 126. The basic framework of multimodal biomedical RAG typically consists of: (1) modality-specific encoders (e.g., Vision Transformers 135 for images, specialized encoders for time-series data 135); (2) knowledge retrieval components that access domain-specific databases based on multimodal query representations; and (3) generation modules that synthesize comprehensive outputs informed by both the multimodal inputs and retrieved knowledge. For instance, Ranjit et al.130, FactMM-RAG 31 and STREAM 133 use multimodal embeddings for retrieval of relevant candidate radiology text to generate Chest X-Ray report. MMed-RAG 131 and RULE 132 build a multimodal RAG system for Medical Visual Question Answering (Medical VQA) and report generation. The MEAG 32 system uses specialized encoders for eye movement data alongside text, Raminedi et al. 129 employ Vision Transformers as encoders for radiological images with cross-attention mechanisms for fusion with text, while Thetbanthad et al. 127 first converts prescription label images to text through OCR before retrieving relevant pharmaceutical information from DrugBank.
@@
Rerankers in biomedical RAG systems serve as intermediary components that refine and prioritize the relevance of initially retrieved documents before they are processed by the generative models 45. Positioned between retrievers and generators, rerankers apply sophisticated algorithms to re-score and re-order candidate documents based on their contextual relevance to the query, effectively functioning as a second-stage filtering mechanism that enhances the precision of information ultimately provided to the generation component. By prioritizing the most pertinent information, rerankers help to minimize hallucinations and improve the veracity of generated outputs.
Biomedical RAG rerankers can be categorized into distinct types based on their underlying methodologies. Statistical rerankers apply traditional information retrieval techniques; for instance, CliniqIR 52 and MEDRAG 101 both implement Reciprocal Rank Fusion (RRF) to combine multiple retrieval signals, while DRAGON-AI 83 employs Maximal Marginal Relevance (MMR) to balance relevance with diversity. Content-based rerankers, in contrast, focus on content similarity; notably, WeiseEule 101 introduces a keyword frequency-based ranking system specifically designed for biomedical literature retrieval. Model-based rerankers employ transformer architectures to perform contextual relevance assessment; specifically, PodGPT 34 uses a general embedding model (bge-reranker-large) for scientific literature retrieval, whereas the CHA2DS2-VASc risk factor extraction system 136 utilizes a specialized cross-encoder model known as BGE M3-Embedding 137 to enhance clinical document relevance scoring. Within this category, some domain adaptation retrievers incorporate a reranker as a component of their model; for example, MedCPT 33 trains a specific reranker denoted as CrossEnc. Additionally, LLM-based rerankers leverage foundation models’ reasoning capabilities; GNQA 77 employs GPT-3.5/4/4o as zero-shot binary classifiers for relevance determination, and LmRaC 138 implements an LLM-based paragraph usefulness assessment with a minimum threshold score of 7. (on a 10 point scale). Finally, hybrid approaches combine multiple strategies to enhance performance. For example, Clinfo.ai 14 combines LLM classification with BM25 to re-rank retrieved articles.
@@
The generation component represents the last element in biomedical RAG systems, responsible for synthesizing retrieved information into human-like responses that address user queries. In biomedical contexts, this process requires balancing natural language fluency with domain-specific accuracy, as generated content often informs healthcare decisions 139. This section examines various generation models employed in recent biomedical RAG studies.

### v3

@@ Introduction
DRAGON-AI toro2024dragonai OpenAI Text-Embedding openai_embeddings with Chroma DB Ontology generation 2024
RISE wang2024diabetesrag OpenAI Text-Embedding with FAISS Diabetes-related inquiries 2024
Dual RAG lee2024dualrag UPstage API and OpenAI Text-Embedding Diabetes management 2024
FAVOR-GPT li2024favorgpt Open AI Text-Embedding with Weaviate DB Genome variant annotations 2024
Endo-chat fu2024application Open AI Text-Embedding with Faiss medical QA for gastrointestinal endoscopy 2024
ChatENT long2024chatent Open AI Text-Embedding Medical QA in otolaryngology 2024
GuideGPT steybe2025evaluation Open AI Text-Embedding MRONJ QA on prevention, diagnosis, and treatment 2024
RECTIFIER unlu2024retrieval Open AI Text-Embedding with Faiss Clinical trial screening for heart failure patients 2024
RAP azimi2025evaluation Amazon Titan Text Embeddings v2 aws_titan_embedding Nutrition-related question answering 2025
InfectA-Chat selcuk2025infecta Open AI Text-Embedding Infectious disease monitoring and QA 2025
DCRAG kainer2025effectiveness OpenAI Text-Embedding Auto-annotation of plant phenotype 2025
@@
PM-Search Jin2022 BioBERT 10.1093/bioinformatics/btz682 Clinical literature retrieval 2022
LADER LADER PubMedBERT pubmedbert Biomedical literature retrieval 2023
CLEAR lopez2025clinical BioBERT Clinical information extraction 2025
KG-RAG Soman2024 PubMedBERT Biomedical Multiple-choice questions 2024
WeiseEule Aftab2024 MedCPT, BioBERT, BioGPT luo2022biogpt with Pinecone DB Biomedical QA 2024
@@
BiomedRAG li2025biomedrag MedLLaMA 13b Biomedical NLP tasks 2024
SurgeryLLM ong2024surgeryllm GPT4All anand2023gpt4all with Chroma DB Surgical decision support 2024
Myers et al. Myers2025 LLM2Vec (for comparison) llm2vec Clinical Information retrieval 2024
@@
Self-BioRAG jeong2024improving MedCPT jin2023medcpt Medical QA 2024
@@
WeiseEule Aftab2024 MedCPT, BioBERT, BioGPT with Pinecone DB Biomedical QA 2024
CliniqIR Abdullahi2024 MedCPT and BM25 Diagnostic decision support 2024
@@
MEDRAG Aftab2024 BM25, SPECTER specter2020cohan, Contriever izacard2021contrieverand MedCPT Biomedical RAG Tool 2024
@@
RAMIE zhan2025ramie MedCPT, Contriever and BMRetriever xu2024bmretriever Biomedical IR about dietary supplements 2025
@@
Type 2: Commercial embedding APIs. The proliferation of foundation models has led to the emergence of proprietary embedding solutions offered by major AI private organizations. These commercial embedding APIs have gained significant adoption in biomedical research, despite their associated costs, due to several compelling advantages: i) they typically leverage substantially larger pretraining datasets, ii) benefit from extensive computational resources during development, and iii) receive continuous refinements based on diverse user interactions. As shown in Table LABEL:tab:Classification, OpenAI’s embedding models dominate this category in recent studies, underscoring their enhanced semantic representation capabilities. While OpenAI’s offerings currently dominate, alternative solutions are also emerging. For example, the RAP system azimi2025evaluation utilizes Amazon’s Titan Text Embeddings v2 aws_titan_embedding for nutrition-related question answering, and other commercial APIs, such as Anthropic’s Voyage-2 anthropic_embeddings and Google’s Vertex AI embeddings google_vertexai_embeddings, though not represented in Table LABEL:tab:Classification, can also be employed in biomedical RAG applications
@@
Type 3: Domain adaptation PLMs. Domain-adapted PLMs are language models that have been specifically pre-trained or fine-tuned on biomedical and clinical corpora, allowing them to develop specialized representations of medical terminology, concepts, and relationships 10.1093/bioinformatics/btz682, pubmedbert. These models mitigate a fundamental challenge in biomedical information retrieval: general-domain language models often struggle with the highly specialized vocabulary, complex semantic relationships, and unique linguistic patterns prevalent in biomedical literature Lyu2024LanguageMA. By adapting pre-training to include PubMed articles, clinical notes, or other domain-specific corpora, these models develop a more nuanced understanding of biomedical language, resulting in competitively improved performance on retrieval tasks involving medical concepts and relationships compared to their general-domain counterparts Lyu2024LanguageMA. As shown in Table LABEL:tab:Classification, examples of these domain-adapted models include PubMedBERT pubmedbert, BioBERT 10.1093/bioinformatics/btz682, and BioGPT luo2022biogpt, which have been widely adopted in biomedical RAG systems. For instance, PM-Search Jin2022 utilized BioBERT for clinical literature retrieval, effectively leveraging the model’s domain-specific representations to improve search precision. Similarly, PubMedBERT has been successfully deployed in LADER LADER for biomedical literature retrieval and in KG-RAG Soman2024 for answering biomedical multiple-choice questions, demonstrating its versatility across different tasks. More recent systems like WeiseEule Aftab2024 compare multiple biomedical models in their study, including BioBERT, BioGPT, and MedCPT jin2023medcpt, to achieve enhanced performance in biomedical QA.
@@
Type 4: LLM-based embedding. Generative LLMs have demonstrated remarkable generative capabilities across diverse NLP tasks, owing to their sophisticated understanding of language and context. With billions of parameters and training on vast corpora, these models develop rich internal representations that capture nuanced semantic relationships and contextual dependencies in text Zhao2023LLMSurvey. Recognizing this powerful language comprehension ability, researchers have begun leveraging open-source LLMs as embedding generators for retrieval tasks, extracting representations from their internal layers to create semantically meaningful document and query embeddings. As shown in Table LABEL:tab:Classification, notable implementations include BiomedRAG li2025biomedrag, which utilizes MedLLaMA-13B wu2023pmc; SurgeryLLM ong2024surgeryllm, which employs GPT4All anand2023gpt4all; and the comparative study by Myers et al. Myers2025, which evaluates LLM2Vec llm2vec for clinical information retrieval applications. However, most LLMs are decoder-only architectures that employ causal attention mechanisms, which inherently limit their ability to generate rich contextualized representations. Thus, researchers are increasingly focusing on developing more advanced LLM-based embedding methods, such as Llama2Vec li-etal-2024-llama2vec, Landmark Embedding luo-etal-2024-landmark, which provide promising dense retrieval alternatives in biomedical RAG.
@@
Type 5: Domain adaptation retrievers. Although the aforementioned off-the-shelf retrievers are readily available, the task of searching for relevant and accurate documents in the medical domain remains challenging. Consequently, customized retrievers tailored specifically to scientific and biomedical tasks have been developed. As shown in Table LABEL:tab:Classification, these advanced retrievers include Contriever izacard2021contriever, MedCPT jin2023medcpt, SPECTER specter2020cohan, BMRetriever xu2024bmretriever, and the MEDRAG toolkit li2025biomedrag. These specialized retrieval architectures demonstrate the importance of domain adaptation in biomedical information retrieval, significantly improving retrieval performance in several studies jeong2024improving, Abdullahi2024, zhan2025ramie. For example, Jeong et al. jeong2024improving proposed Self-BioRAG, which leverages MedCPT to improve medical question answering through a self-augmentation approach. Similarly, the MEDRAG framework Aftab2024 combines multiple retrieval strategies, including BM25, SPECTER, Contriever, and MedCPT to create a comprehensive biomedical RAG tool. RAMIEzhan2025ramie utilizes MedCPT, Contriever, and BMRetriever as retrievers, builds a RAG framework to extract diverse types of information about dietary supplements (DSs) from clinical records, enhanced overall accuracy in its experiments zhan2025ramie. These implementations highlight how domain-adapted retrievers benefit biomedical RAG systems.
@@
Chroma chroma2023 AI-native open-source vector database with built-in functionality
Faiss faiss2023 Library for efficient similarity search and clustering of dense vectors
Pinecone pinecone_overview Production-ready vector database for similarity search at scale
Weaviate weaviate_developers Graph-based vector search engine with semantic capabilities
HNSW malkov2016efficient Hierarchical navigable small world graphs for approximate nearest neighbor search
Dense Retriever Infrastructure. Vector stores provide an infrastructure element for dense retrievers in semantic search systems han2023comprehensive. These specialized databases efficiently index, store, and search high-dimensional vector embeddings, offering optimized methods like K-nearest neighbors (KNN) for RAG developers. Vector databases enable rapid retrieval of semantically similar documents at scale, thereby improving the efficiency and reducing the response time of biomedical RAG systems. The process of embedding, storing, and searching documents involves transforming text into vector representations, organizing them in specialized data structures, and efficiently retrieving the most relevant matches. Based on our survey of biomedical RAG publications presented in Table LABEL:tab:Classification, we compiled the most commonly used vector databases in Table 4, among which FAISS and Chroma emerge as the most popular choices, being utilized 4 and 3 times, respectively.
@@
Knowledge graph retrievers provide structured, interlinked domain knowledge in biomedical RAG, helping to address several limitations of traditional retrieval methods. Unlike vector-based approaches, knowledge graphs capture explicit relationships between biomedical entities, enabling more precise retrieval of contextually relevant information, enhanced reasoning capabilities, and improved explainabilitypeng2024graph. Table 5 presents recent biomedical RAG studies that incorporate knowledge graphs, outlining their application domains and the knowledge graph sources they employ. The application areas covered by these systems include medical question answering Soman2024, knowledge extraction DiMaria2024, and diagnosis prediction Gao2025.
@@
KG-RAG Soman2024 Biomedical text generation, medical question answering SPOKE biomedical knowledge graph 2024
HEALIE Kakalou2024 Personalized medical content generation HEALIE KG (customized) 2024
KRAGEN Matsumoto2024 Biomedical problem solving Alzheimer’s knowledge graph (AlzKB) 2024
Ascle Ascle Various medical text generation tasks UMLS 2024
DALK li2024dalk Alzheimer’s QA Alzheimer’s knowledge graph (AlzKB) 2024
Gilbert et al. Gilbert2024 Medical information structuring and interlinking Not specified 2024
NetMe 2.0 DiMaria2024 Biomedical knowledge extraction BKGs constructed using OntoTagMe and Wikidata 2024
NEKO XIAO202560 Knowledge mining in synthetic biology PubMed-derived knowledge graphs 2025
DR.KNOWS Gao2025 Diagnosis prediction from EHRs UMLS 2025
ESCARGOT Matsumoto2025 Biomedical reasoning and knowledge retrieval Alzheimer’s knowledge graph (AlzKB) 2025
Figure 3: The Framework of Knowledge Graphs-Based Biomedical RAG
To further clarify how knowledge graphs are incorporated into biomedical RAG systems, Fig. 3 illustrates the integration of knowledge graphs with LLMs through four key components:
Biomedical knowledge graph construction or curation involves the systematic process of building, updating, or adapting knowledge graphs from domain-specific ontologies, literature, and other data sources to create structured biomedical knowledge repositories. For example, NetMe 2.0 DiMaria2024 constructs biomedical knowledge graphs on-the-fly using OntoTagMe for entity linking, while NEKO XIAO202560 mines PubMed literature to build knowledge graphs for synthetic biology. By contrast, KG-RAG Soman2024 utilizes the pre-existing SPOKE biomedical knowledge graph.
Query structuration encompasses the conversion of natural language queries or user inputs into structured queries that efficiently retrieve relevant subgraphs from the knowledge graph. For instance, ESCARGOT Matsumoto2025 employs Cypher queries for precise retrieval, KRAGEN Matsumoto2024 vectorizes knowledge graphs for semantic searching, and DR.KNOWS Gao2025 transforms queries to retrieve UMLS-based diagnostic pathways aligned with patient data.
Path-based or entity-centric knowledge extraction refers to the process of extracting relevant knowledge patterns, paths, or entity-centered information from the knowledge graph based on the transformed query. Specifically, DR.KNOWS Gao2025 utilizes attention-based path ranking for diagnostic reasoning, KRAGEN Matsumoto2024 applies graph-of-thoughts for multi-hop relationships, and HEALIE Kakalou2024 extracts paths incorporating both medical knowledge and socio-economic factors.
Strategic integration into LLM prompts involves methods for incorporating retrieved knowledge into prompts in ways that maximize factual accuracy while optimizing token usage and computational efficiency. For example, KG-RAG Soman2024 implements token-optimized retrieval to reduce consumption, whereas ESCARGOT Matsumoto2025 selectively integrates knowledge through a dynamic Graph of Thoughts. In addition, NetMe 2.0 DiMaria2024 transforms graph data into natural language prompts via its Graph-RAG module.
@@
Multimodal biomedical RAG represents an advanced approach to healthcare AI that integrates LLMs with multiple types of data inputs beyond text, such as medical images, electrocardiogram (ECG), and other clinical measurements. These systems can enhance diagnostic accuracy and clinical decision support by leveraging the complementary strengths of different data modalities, retrieving relevant information from specialized knowledge bases, and generating comprehensive insights that would be impossible to achieve with single-modality approaches abootorabi2025askmodalitycomprehensivesurvey. As shown in Table 6, multimodal biomedical RAG systems have been successfully deployed for amblyopia detection Upadhyaya2024, prescription label interpretation Thetbanthad2025, lung cancer staging Tozuka2024, digital pathology analysis Omar2024, and automated radiology report generation pmlr-v219-ranjit23a, Sun2024FactAware, xia2024mmedrag, Xia2024RULE, Yang2025STREAM.
@@
Ranjit et al. pmlr-v219-ranjit23a Chest X-Ray report generation Chest X-ray and text 2023
FactMM-RAG Sun2024FactAware Radiology report generation X-Ray and text 2024
MMed-RAG xia2024mmedrag Medical VQA & Report generation Medical images and text 2024
MEAG Upadhyaya2024 Amblyopia diagnosis Eye tracking and text 2024
Tozuka et al. Tozuka2024 Lung cancer staging CT findings and TNM classifications 2024
GPT4DFCI-RAG Omar2024 Digital pathology QA Pathology images and text 2024
Raminedi et al. Raminedi2024MultimodalTransformer Radiology report generation X-ray images and text 2024
RULE Xia2024RULE Medical VQA & Report generation Medical images and text 2024
Thetbanthad et al. Thetbanthad2025 Prescription label identification Image (labels) and text 2025
Hu et al. Hu2025 Pathology report generation Whole slide images and text 2025
STREAM Yang2025STREAM Chest X-ray report generation Chest X-ray and text 2025
Figure 4 illustrates the architectural workflow of a multimodal biomedical RAG system. The key distinction between multimodal biomedical RAG and traditional biomedical RAG lies in data input diversity and processing architecture. While traditional RAG systems primarily process textual queries to retrieve text-based medical knowledge, multimodal RAG incorporates and processes non-textual data (images, sensor readings, etc.) alongside text, requiring specialized encoders for each modality to meaningfully integrate these diverse inputs abootorabi2025askmodalitycomprehensivesurvey. The basic framework of multimodal biomedical RAG typically consists of: (1) modality-specific encoders (e.g., Vision Transformers chen2021outperform for images, specialized encoders for time-series data chen2021outperform); (2) knowledge retrieval components that access domain-specific databases based on multimodal query representations; and (3) generation modules that synthesize comprehensive outputs informed by both the multimodal inputs and retrieved knowledge. For instance, Ranjit et al.pmlr-v219-ranjit23a, FactMM-RAG Sun2024FactAware and STREAM Yang2025STREAM use multimodal embeddings for retrieval of relevant candidate radiology text to generate Chest X-Ray report. MMed-RAG xia2024mmedrag and RULE Xia2024RULE build a multimodal RAG system for Medical Visual Question Answering (Medical VQA) and report generation. The MEAG Upadhyaya2024 system uses specialized encoders for eye movement data alongside text, Raminedi et al. Omar2024 employ Vision Transformers as encoders for radiological images with cross-attention mechanisms for fusion with text, while Thetbanthad et al. Thetbanthad2025 first converts prescription label images to text through OCR before retrieving relevant pharmaceutical information from DrugBank.
@@
Rerankers in biomedical RAG systems serve as intermediary components that refine and prioritize the relevance of initially retrieved documents before they are processed by the generative models sivarajkumar2024clinical. Positioned between retrievers and generators, rerankers apply appropriate ranking algorithms to re-score and re-order candidate documents based on their contextual relevance to the query, effectively functioning as a second-stage filtering mechanism that enhances the precision of information ultimately provided to the generation component. By prioritizing the most pertinent information, rerankers help to minimize hallucinations and improve the veracity of generated outputs.
Biomedical RAG rerankers can be categorized into distinct types based on their underlying methodologies. Statistical rerankers apply traditional information retrieval techniques; for instance, CliniqIR Abdullahi2024 and MEDRAG Aftab2024 both implement Reciprocal Rank Fusion (RRF) to combine multiple retrieval signals, while DRAGON-AI toro2024dragonai employs Maximal Marginal Relevance (MMR) to balance relevance with diversity. In addition, WeiseEule Aftab2024 introduces a keyword frequency–based ranking system tailored for biomedical literature retrieval. Model-based rerankers employ transformer architectures to perform contextual relevance assessment; specifically, PodGPT jia2024podgpt uses a general embedding model (bge-reranker-large) for scientific literature retrieval, whereas the CHA2DS2-VASc risk factor extraction system adejumo2024retrieval utilizes a specialized cross-encoder model known as BGE M3-Embedding chen2024m3embedding to enhance clinical document relevance scoring. Within this category, some domain adaptation retrievers incorporate a reranker as a component of their model; for example, MedCPT jin2023medcpt trains a specific reranker denoted as CrossEnc. Additionally, LLM-based rerankers leverage foundation models’ reasoning capabilities; GNQA darnell2024biomedical employs GPT-3.5/4/4o as zero-shot binary classifiers for relevance determination, and LmRaC craig2024lmrac implements an LLM-based paragraph usefulness assessment with a minimum threshold score of 7. (on a 10 point scale). Finally, hybrid approaches combine multiple strategies to enhance performance. For example, Clinfo.ai Lozano2024ClinfoAI combines LLM classification with BM25 to re-rank retrieved articles.
@@
The generation component represents the last element in biomedical RAG systems, responsible for synthesizing retrieved information into human-like responses that address user queries. In biomedical contexts, this process requires balancing natural language fluency with domain-specific accuracy, as generated content often informs healthcare decisions lee2024enhancing. This section examines various generation models employed in recent biomedical RAG studies.